OpenStudy (anonymous):

Here is the question, and i think I have the equation set up correctly, I just need to know how to actually solve it. :/ Natasha rides her bike (at a constant speed) for 4 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 10 hours. Find find the total round trip distance she traveled. (s + 3)4 = d/2 (s - 3)10 = d/2

OpenStudy (anonymous):

1º step towards the solution:) (s + 3)4 = (s - 3)10

OpenStudy (blank ):

The correct way of writing @myko is: (s + 3)(4) = (s - 3)(10) The result is: s=7

OpenStudy (blank ):

(7 + 3)(4) = 40 (7 - 3)(10) = 40 40 + 40 = 80

OpenStudy (blank ):

Distance is 80 miles.

OpenStudy (anonymous):

Thanks, to both of you. That's what I thought but was second guessing myself too much. It's been a while since I've had to do this.

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

ha!